### v3

# Few-Shot Learning on Graphs: from Meta-learning to Pre-training and Prompting

Abstract
Graph representation learning, a critical step in graph-centric tasks, has seen significant advancements. Earlier techniques often operate in an end-to-end setting, where performance heavily relies on the availability of ample labeled data. This constraint has spurred the emergence of few-shot learning on graphs, where only a few task-specific labels are available for each task. Given the extensive literature in this field, this survey endeavors to synthesize recent developments, provide comparative insights, and identify future directions. We systematically categorize existing studies into three major families: meta-learning approaches, pre-training approaches, and hybrid approaches, with a finer-grained classification in each family to aid readers in their method selection process. Within each category, we analyze the relationships among these methods and compare their strengths and limitations. Finally, we outline prospective future directions for few-shot learning on graphs to catalyze continued innovation in this field.
1 Introduction
Data objects are often involved in complex interactions with each other, forming diverse network structures also known as graphs. The prevalence of graph-structured data gives rise to a surge in graph-centric tasks, such as node classification, link prediction, and graph classification. Recently, graph representation learning in a supervised setting Kipf and Welling (2017) has emerged as a prominent family of techniques for these graph-based tasks. The success of such techniques often depend on extensive labeled data for end-to-end training. However, acquiring labeled data is often challenging or expensive, limiting the performance of these supervised techniques Liu et al. (2023c).
Figure 1: Taxonomy of techniques for few-shot learning on graphs.
In light of this issue, few-shot learning on graphs Wang et al. (2020) has attracted significant interest in the field. Specifically, the number of labeled instances (termed as shots) in each task is notably small, representing a more practical setting in real-world applications. Few-shot learning approaches have been widely adopted in graph-based tasks, ranging from social network analysis to recommendation systems and molecular research. These approaches boil down to learning prior insights from data, often external to the target tasks, and subsequently adapting the prior knowledge to the downstream target tasks in a few-shot setup. Naturally, two key challenges arise. First, how can we learn a high-quality prior from graph data to provide useful references for downstream few-shot tasks? Second, how can we effectively and efficiently adapt the prior to downstream tasks?
To answer these challenges, this survey presents a concise yet systematic review of prevailing methods for few-shot learning on graphs. It aims to summarize recent developments, provide comparative insights, and identify future research avenues, contributing substantively to the few-shot learning and graph learning communities. For a systematic and comprehensive overview, we categorize the body of literature for few-shot learning on graphs into three major families: (1) Meta-learning approaches, (2) Pre-training approaches, and (3) Hybrid approaches, as shown in Fig. 1. Notably, both meta-learning Ding et al. (2020); Sheng et al. (2020) and pre-training Veličković et al. (2018); Qiu et al. (2020) represent two predominant research trends to learn priors from additional data, which are subsequently adapted for few-shot downstream tasks. Nevertheless, the two trends embrace distinct paradigms. On one hand, mainstream meta-learning approaches, such as Model Agnostic Meta-Learning (MAML) Finn et al. (2017) and Prototypical Networks (Protonets) Snell et al. (2017), derive a prior from a series of “meta-training” tasks (with task-specific labels) that mirror the downstream “meta-testing” tasks, assuming that all these tasks are drawn from an identical task distribution. They do not directly learn a model to serve as the prior. Rather, they aspire to learn a general adaptation process, also known as “learning-to-learn.” On the other hand, pre-training approaches utilize unlabeled data to optimize self-supervised pretext tasks, generally unrelated to the downstream tasks. They directly train a model as the prior, followed by an additional tuning phase to adapt the prior to specific downstream tasks. In the context of few-shot learning on graphs, meta-learning methods further exploit structure- or adaptation-based enhancements to cater to graph data. In contrast, pre-training approaches commonly train a graph encoder (i.e., the prior) in the pre-training phase, aiming to capture inherent characteristics of graph data, such as local connectivity or global topological patterns. Then, in the downstream phase, they either perform a full fine-tuning of the pre-trained graph encoder on each task, or employ a more parameter-efficient adaptation strategy. Finally, hybrid approaches integrating both paradigms have emerged recently, thereby synergistically harnessing the strengths of meta-learning and pre-training approaches. That is, they are able to exploit both unlabeled graph data and annotations on similar tasks, for a more seamless adaptation to downstream tasks.
Over the past few years, many efforts have been devoted to reviewing the few-shot learning literature from diverse perspectives Wang et al. (2020); Song et al. (2023); Wang et al. (2019a), but they do not cover few-shot learning on graphs. Recently, Zhang et al. Zhang et al. (2022) have discussed this field but limited its categorization to node-, edge- and graph-level problems in the meta-learning paradigm, lacking a comprehensive view of the technical approaches. Our survey endeavors to fill this gap by presenting a holistic overview of few-shot learning on graphs, proposing a fine-grained taxonomy of both classical and emerging methodologies. In summary, the contribution of our survey is threefold: 1) introduction of a novel taxonomy, enabling a systematic and concise overview of the literature in the field; 2) synthesis of recent progress within each branch of the taxonomy, coupled with comparative analyses of their respective strengths and weaknesses; 3) discussion of open challenges and promising directions for future research.
The remainder of this survey is organized as follows. Sects. 2 and 3 delve into the meta-learning and pre-training paradigms, respectively, in the context of few-shot learning on graphs. Sect. 4 explores the hybrid approaches that integrate and benefit from both paradigms. Sect. 5 examines open challenges in the field and identifies potential avenues for future research. Finally, Sect. 6 concludes the survey.
2 Meta-learning Approaches
Figure 2: An illustration of meta-learning on graph.
Prevailing episodic meta-learning methods are designed to learn a prior from a series of meta-training tasks, which is transferable to the downstream meta-testing tasks. The concept is illustrated in Fig. 2, using few-shot node classification on a graph as an example. In this scenario, meta-training tasks are constructed using base classes with ample labels (“base set”), whereas a meta-testing task involves novel classes with only few-shot labels. Each task comprises a support set and a query set, which function as training and testing data, respectively, within each task Finn et al. (2017). The support set is always labeled, whereas the query set is only labeled in meta-training but is unlabeled for inference and prediction in meta-testing. The prior is optimized on the meta-training tasks to enable rapid, lightweight adaptation using the support set that can minimize the loss on the query set. Hence, for the prior to be transferable to meta-testing tasks, such meta-learners assume that the meta-training and -testing tasks are sampled from an i.i.d. task distribution.
To tailor to few-shot learning on graphs, standard meta-learning methodologies have been enhanced in various ways. We further categorize the literature according to their graph-specific enhancements within the meta-learning framework, namely, structure-based enhancement and adaptation-based enhancement, as summarized in Table 1.
Method Enhancement Meta-learner Task
Node Edge Graph
GPN Ding et al. (2020) structure (node) Protonets $\checkmark$ $\times$ $\times$
FAAN Sheng et al. (2020) structure (node) Protonets $\times$ $\checkmark$ $\times$
HMNet Xiao et al. (2021) structure (edge) MN $\times$ $\checkmark$ $\times$
Wang et al. (2019b) structure (edge) MAML $\times$ $\checkmark$ $\times$
RALE Liu et al. (2021b) structure (path) MAML $\checkmark$ $\times$ $\times$
G-Meta Huang and Zitnik (2020) structure (subgraph) hybrid $\checkmark$ $\checkmark$ $\times$
GEN Baek et al. (2020) structure (subgraph) Protonets $\times$ $\checkmark$ $\times$
GFL Yao et al. (2020) adaptation (graph) MAML $\checkmark$ $\times$ $\times$
MI-GNN Wen et al. (2021) adaptation (graph) hybrid $\checkmark$ $\times$ $\times$
MetaTNE Snell et al. (2017) adaptation (task) Protonets $\checkmark$ $\times$ $\times$
AS-MAML Ma et al. (2020) adaptation (step size) MAML $\times$ $\times$ $\checkmark$
Table 1: Meta-learning approaches for few-shot learning on graphs.
2.1 Structure-based enhancement
A unique characteristic of graph data lies in the connectivity structures between nodes, presenting crucial information for coping with graph data. Consequently, many meta-learning methods for graph-based tasks have focused on exploiting graph structures to enhance the learning of the prior. Based on the type of graph structures, they can be further grouped into node-, edge-, and subgraph-level enhancements.
Node-level enhancement. The nodes of a graph encompass implicit structural information for enhancing the meta-learning model. Recent studies predominantly focus on differentiating node weights in a task to reflect their varying structural importance. Specifically, in Protonets-based methods such as GPN Ding et al. (2020) and FAAN Sheng et al. (2020), it is observed that each support node contributes variably to the prototype of each class in a task. Thus, they propose to evaluate the importance of nodes in the support set and adjust their contribution w.r.t. the query node when constructing the prototypes.
Edge-level enhancement. Edges are fundamental to graphs as they represent explicit first-order structures. Moreover, a sequence of edges forms a path, which is capable of representing more complex high-order interactions among distant nodes. As a result, many meta-learning methods on graphs incorporate edge- or path-based enhancements. On one hand, edge-based enhancement leverages auxiliary information associated with edges, thereby extending beyond mere node-level information. For instance, to boost few-shot link prediction in knowledge graphs, HMNet Xiao et al. (2021) adopts matching networks (MN) as its meta-learning framework, performing matching for both entities and relations. In particular, for relation-level matching, it calculates a relation-aware score using a relation encoder to enrich the link prediction task. Another work Wang et al. (2019b) develops a generative model to generate new triplets, serving to augment the sparse links within a MAML-based framework. On the other hand, paths can capture long-range dependencies between distant nodes. RALE Liu et al. (2021b) proposes to exploit the paths between each query node and the support nodes for task-level dependencies, and between each target node and the hub nodes for graph-level dependencies.
Subgraph-level enhancement. Subgraphs provide rich contextual information to its containing nodes. Hence, recent studies have utilized subgraphs to enhance the expressiveness of nodes. G-Meta Huang and Zitnik (2020) generates class prototypes based on the contextual subgraph of each support node, and similarly expands each query node to its subgraph, to enhance the robustness of prototype and query representations. Similarly, GEN Baek et al. (2020) proposes a graph extrapolation network to extrapolate knowledge through the neighbors of the support set, which essentially exploits the one-hop subgraphs of the support nodes.
2.2 Adaptation-based enhancement
Achieving rapid adaptation to the query set through the support set is a key requirement of meta-learning. In particular, adaptation enables the customization of a globally shared prior into a localized or specialized model for each task, thereby not only preserving the unique traits of each individual task but also capitalizing on shared common knowledge. Hence, many approaches aim to enhance the adaptation mechanism for few-shot learning on graphs.
GFL Yao et al. (2020) recognizes the topological variances across different graphs, and thus customizes a global prior for each individual graph. More concretely, it generates class prototypes specifically tailored to each graph, utilizing a gate function based on graph-specific representation that is applied to the global prior. MI-GNN Wen et al. (2021) adopts a similar idea and employs a hypernetworks Perez et al. (2018) to modulate the global prior for each graph before undergoing a further task-wise adaptation within a MAML framework. Meanwhile, there is potential for enhancement in task-wise adaptation. In the context of multi-label few-shot classification, MetaTNE Lan et al. (2020) observes an increase in task diversity due to the multi-label setting. Thus, adopting a uniform node embedding across such diverse tasks may result in inadequate expressiveness, given that the same node could be associated with different labels in different tasks. To address this, MetaTNE conducts an additional adaptation for the node embeddings w.r.t. the query set in each task, thereby better adapting to the characteristics of the task. In addition to enhancing adaptation from the perspective of graph or task, AS-MAML Ma et al. (2020) improves adaptation from an optimization standpoint. It employs a reinforcement learning-based controller to determine the optimal step size for the adaptation process in MAML.
2.3 Discussion
In summary, existing research for few-shot learning on graphs often enhances a standard meta-learner, either through structural augmentation or by refining the adaptation process. However, these approaches share a common drawback inherent to meta-learning approaches. First, they still require a large volume of labels for a “base set” during the meta-training phase, which can be challenging or costly to acquire. Second, they fail to leverage the vast amount of unlabeled data to learn a more comprehensive prior. Third, they are limited by the assumption in task distribution, and cannot handle different types of tasks downstream. This leads to a further challenge: How can we address a diverse range of few-shot learning tasks on graphs without an extensively annotated base set, while utilizing abundant unlabeled graph data? Recently, pre-training on graphs has emerged as a promising solution to this challenge.
3 Pre-training Approaches
Pre-training for few-shot learning on graphs typically utilizes self-supervised methods on unlabeled graph data to train a graph encoder, which aims to capture task-agnostic intrinsic knowledge in graphs, in the pre-training stage. Typical knowledge include as local and global structures, as well as node features. The pre-trained graph encoder, serving as a form of prior knowledge, can be used to address a variety of downstream tasks through an adaptation stage. An overview of pre-training and adaptation process is depicted in Fig. 3.
Given the abundance of unlabeled graph data in many domains and the flexibility to deal with many types of downstream tasks, pre-training approaches have been popularized. In this section, we provide a taxonomy of existing work in this direction, covering both the pre-training stage and the downstream adaptation stage.
Figure 3: An illustration of pre-training and adaptation for few-shot learning on graphs.
Method Target Pos/Neg samples Augmentation
instance
GRACE Zhu et al. (2020) node node/node random
GCC Qiu et al. (2020) subgraph subgraph/subgraph random
GraphCL You et al. (2020) graph graph/graph random
DGI Veličković et al. (2018) graph node/graph random
InfoGraph Sun et al. (2019) graph node/graph random
MVGRL Hassani and Khasahmadi (2020) subgraph node/subgraph diffusion
JOAO You et al. (2021) graph graph/graph automated
Table 2: Summary of contrastive pre-training strategies.
3.1 Strategies for the pre-training stage
Current self-supervised tasks (also known as pretext tasks) for graph pre-training predominantly fall into two broad categories: contrastive and generative strategies.
Contrastive strategies. Contrastive pre-training functions by contrasting instances at various scales within a graph. For each target instance, contrastive methods sample positive and negative instances, aiming to bring the positive instances closer to the target in the latent space while distancing the negative instances. A summary of contrastive strategies is presented in Table 2.
First, the scale of instances used for contrasting varies, ranging from individual nodes to broader subgraphs, or even the entire graph. One one hand, node-level approaches provide a more granular treatment focusing on the representations of each individual node. An example is GRACE Zhu et al. (2020), which generates two augmented views by randomly corrupting the original graph. Then, for each target node in one view, it treats the corresponding node from the other view as positive, while others as negative. On the other hand, graph-level approaches expand the scope to (sub)graphs, thereby capturing global structural patterns. GCC Qiu et al. (2020) conducts random walk in an $r$-hop ego network to obtain a target subgraph instance alongside its positive instances. Negative instances are acquired by performing random walk in an $r^{\prime}$-hop ego network of the same ego node, for some $r\neq r^{\prime}$. Similarly, GraphCL You et al. (2020) generates graph views with various data augmentations and treats graph views augmented from the same graph as positive instances, while those from different graphs as negative instances. Furthermore, there exist cross-scale approaches contrasting between node- and graph-level instances, such as DGI Veličković et al. (2018) and InfoGraph Sun et al. (2019). They both maximize the mutual information between the local node representation and the global graph representation, but differ in their way of generating negative instances.
Second, to sample the positive or negative instances, different augmentation techniques have been proposed for learning robust representations. The most straightforward augmentation strategy perturbs or samples the graphs in a uniformly random fashion Veličković et al. (2018); Zhu et al. (2020), while recent studies turn to a more strategic selection of augmentations. On one hand, MVGRL Hassani and Khasahmadi (2020) applies graph diffusion to transform the adjacency matrix as structure-space augmentation. On the other hand, JOAO You et al. (2021) presents an automated framework that exploits the most challenging augmentation relative to the current loss.
Method Reconstruction objective
node feature node degree edge adj. matrix graph feature
VGAE Kipf and Welling (2016) $\times$ $\times$ $\times$ $\checkmark$ $\times$
GPT-GNN Hu et al. (2020b) $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$
LaGraph Xie et al. (2022) $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$
GraphMAE Hou et al. (2022) $\checkmark$ $\times$ $\times$ $\times$ $\times$
GraphMAE2 Hou et al. (2023) $\checkmark$ $\times$ $\times$ $\times$ $\times$
Table 3: Summary of generative pre-training strategies.
Generative strategies. Concurrent to the development of contrastive strategies, generative methods present a different perspective for graph pre-training. These methods aim to reconstruct parts of the graph, unlike contrastive methods that rely on discriminating different instances on the graph. Various reconstruction objectives have been proposed, including the adjacency matrix Kipf and Welling (2016), masked edges Hu et al. (2020b) or node features Hou et al. (2022). These objectives involve structure reconstruction or feature reconstruction, with many studies employing objectives from both categories, as summarized in Table 3.
Structure reconstruction focuses on the graph topology, aiming to regenerate the entire graph structure or part of it, compelling the graph encoder to preserve the inherent structural patterns. For instance, VGAE Kipf and Welling (2016) attempt to reconstruct the adjacency matrix using the inner product of node embeddings. However, it mainly captures one-hop connectivity patterns, overlooking complex high-order structural information. Alternatively, GPT-GNN Hu et al. (2020b) choose to mask a portion of the edges, and aim to reconstruct these missing edges using the remaining unmasked graph structure. Notably, GPT-GNN pursues a dual goal, which includes the generation of node features. The additional objective helps the encoder recognize the dependency between features and structures, potentially facilitating the modeling of higher-order graph structures.
Feature reconstruction shifts the focus to generating node features in either the original feature space or the latent space. LaGraph Xie et al. (2022) introduces a predictive framework to perform unobserved latent graph prediction, employing node masking to compute the invariance term in its predictive objective. Meanwhile, GraphMAE Hou et al. (2022) turns to masked node feature reconstruction, proposing a masking strategy and error metric designed for robust training. Building upon GraphMAE, GraphMAE2 Hou et al. (2023) introduces two decoding strategies, namely, multi-view random re-mask decoding and latent representation prediction, to regularize feature reconstruction and minimize overfitting.
3.2 Downstream adaptation by fine-tuning
The pre-trained graph encoder entails prior knowledge about the intrinsic properties of the graphs used in the pre-training stage. This prior can be transferred to various downstream tasks by initializing a downstream model with the pre-trained weights. To tailor the initial model to each specific task, it is equipped with a task-specific projection head, and undergoes an adaptation stage known as fine-tuning. During fine-tuning, for each task, the model is further trained with task-specific annotations, updating all parameters in both the pre-trained graph encoder and the task head in the process Qiu et al. (2020); Hu et al. (2020b), as shown in Fig. 3(c), left.
However, adaptation by fine-tuning, while intuitive and straightforward, has its limitations Liu et al. (2023a). First, an objective gap often exists between the pretext tasks in pre-training and the actual downstream tasks, which can compromise the effectiveness of the adaptation process. Second, updating all parameters can be inefficient, especially in the case of large pre-trained models. In our context of few-shot learning, fine-tuning may lead to overfitting due to the limited number of task-specific labels Liu et al. (2023c).
3.3 Parameter-efficient downstream adaptation
To mitigate the problems of adaptation by fine-tuning, recent efforts have focused on parameter-efficient adaptation, which avoids updating all parameters in the pre-trained graph encoder. Key strategies include prompt tuning and parameter-efficient fine-tuning.
Paper Template Feature Structure Multiple Prompt Downstream Task
prompt prompt pretext tasks Initialization Node Edge Graph
GPPT Sun et al. (2022) token-subgraph similarity calculation: $\text{sim}(\mathbf{s}_{v},\vec{t}_{y})$ input $\times$ $\times$ random $\checkmark$ $\times$ $\times$
GraphPrompt Liu et al. (2023c) subgraph similarity calculation: $\text{sim}(\mathbf{s}_{u},\mathbf{s}_{v})$ readout $\times$ $\times$ random $\checkmark$ $\checkmark$ $\checkmark$
GraphPrompt+ Yu et al. (2023b) subgraph similarity calculation: $\text{sim}(\mathbf{s}_{u},\mathbf{s}_{v})$ all layers $\times$ $\times$ random $\checkmark$ $\checkmark$ $\checkmark$
GPF Fang et al. (2023) universal feature space (implicit template) input $\times$ $\times$ random $\checkmark$ $\checkmark$ $\checkmark$
MultiGPrompt Yu et al. (2023c) node similarity calculation: $\text{sim}(\mathbf{h}_{u},\mathbf{h}_{v})$ all layers $\times$ $\checkmark$ pre-trained tokens $\checkmark$ $\checkmark$ $\checkmark$
SGL Zhu et al. (2023) dual-template: $\text{CL}(\mathbf{h}_{u},\mathbf{h}_{v}),\text{GL}(x_{v},\tilde{x}_{v})$$\times$ $\checkmark$ $\checkmark$ random $\times$ $\times$ $\checkmark$
SAP Ge et al. (2023) dual-view similarity: $\text{sim}(\text{MLP}(X),\text{GNN}(X,A))$$\times$ $\checkmark$ $\checkmark$ random $\checkmark$ $\times$ $\checkmark$
ULTRA-DP Chen et al. (2023) node-node / node-group similarity calculation: $\text{sim}(\mathbf{h}_{u},\mathbf{h}_{v})$ input $\checkmark$ $\checkmark$ random $\checkmark$ $\checkmark$ $\times$
HGPrompt Yu et al. (2023a) task template: $\text{sim}(\mathbf{s}_{u},\mathbf{s}_{v})$ graph template: $HG\to\{G^{0},G^{1},G^{2},\ldots\}$ readout $\times$ $\times$ random $\checkmark$ $\checkmark$ $\checkmark$
VNT Tan et al. (2023) node attribute reconstruction: $\text{MSE}(x_{v},\tilde{x}_{v})$ structure recovery: $\text{MSE}(\{\mathbf{h}_{u},\mathbf{h}_{v}\})$ input $\times$ $\times$ meta-trained $\checkmark$ $\times$ $\times$
ProG Sun et al. (2023) subgraph classification: $\text{CLS}(\mathbf{s})$$\times$ $\checkmark$ $\times$ meta-trained $\checkmark$ $\checkmark$ $\checkmark$
Table 4: Summary of prompt tuning on graphs. For ease of comparison, we adapt the notations used in the original references with a consistent set of symbols. $v,u$: Nodes $v,u$; $\vec{t}_{y}$: Prototype token $\mathbf{t}$ of class $y$; $\vec{h}_{v}$: Node $v$’s embedding; $x_{v}$: Node $v$’s attribute; $\mathbf{s}_{v}$: Subgraph of $v$’s embedding; CL: Contrastive learning; GL: Generative learning; CLS: Classification; $G$: Homogeneous graph; $HG$: Heterogeneous graph.
Prompting on graphs. Originating in the field of natural language processing (NLP), prompting has demonstrated its efficacy in adapting pre-trained language models to diverse language tasks Liu et al. (2023a). In essence, a prompt vector is introduced to modify or reformulate the original input for the pre-trained model, drawing the downstream task closer to the pretext tasks under a unified template. Notably, the prompt vector has a negligible size relative to the pre-trained parameters. During the adaptation process, only a very small prompt vector is tuned without updating the pre-trained model, making the approach parameter efficient. Recently, prompt tuning have been extended to graph learning Sun et al. (2022); Liu et al. (2023c). Building on a unified template that aligns the pretext and downstream losses, these methods design prompts to further narrow the gap between pretext and downstream tasks, as illustrated in Fig. 3(c), right. Specific approaches vary in the choice of the template and prompt design, as summarized in Table 4.
In terms of template, similarity calculation has been widely adopted Sun et al. (2022); Liu et al. (2023c); Yu et al. (2023b, c); Chen et al. (2023); Yu et al. (2023a) to unify three major tasks: node classification (NC), link prediction (LP), and graph classification (GC). For instance, GraphPrompt Liu et al. (2023c) reformulates LP as calculating the similarity between the ego subgraphs of two nodes. Then, it applies the same template to NC and GC, treating them as calculating the similarity between the subgraph of the target instance (using an ego network for NC, or the target graph itself for GC) and the prototypical subgraph for each class. While GraphPrompt only adopts LP as the pretext task, GraphPrompt+ Yu et al. (2023b) illustrates that other prevailing contrastive pretext tasks can also be redefined using subgraph similarity. Likewise, SGL Zhu et al. (2023) and SAP Ge et al. (2023) also employ a form of similarity calculation for contrastive tasks. SGL further incorporates a generative pretext task based on mask node feature reconstruction, which is employed as an additional template for downstream prompt learning. In contrast, GPF Fang et al. (2023) does not employ an explicit template to align task formats. Instead, it assumes a universal feature space as an implicit template to unify different tasks.
In terms of prompt design, it can be further categorized into feature prompt and structure prompt. On one hand, similar to prompt in NLP Liu et al. (2023a), feature prompt modifies node features, reformulating the original downstream tasks to narrow the gap from the pre-trained model. Some studies incorporate feature prompts in the input layer Sun et al. (2022); Fang et al. (2023); Chen et al. (2023) to explicit modify the input in a task-specific manner, while others apply them to the readout layer to guide subgraph readout Liu et al. (2023c); Yu et al. (2023a), or all layers of the pre-trained model Yu et al. (2023b, c). On the other hand, a structure prompt acts as a virtual graph structure to reformulate the implicit structural input for downstream tasks. Structure prompts can take the form of virtual nodes or graphs representing class prototypes for classification tasks Zhu et al. (2023); Ge et al. (2023); Chen et al. (2023), or virtual nodes that augment link prediction Chen et al. (2023).
To learn a comprehensive range of prior knowledge from diverse perspectives, some studies propose to utilize multiple pretext tasks Zhu et al. (2023); Ge et al. (2023); Chen et al. (2023). However, directly combining various pretext tasks may result in task interference Yu et al. (2023c), compromising the complementary nature of employing multiple tasks. To address this issue, MultiGPrompt Yu et al. (2023c) utilizes a series of pretext tokens to facilitate cooperation among pretext tasks. It further proposes a composed prompt, which is initialized based on the pre-trained pretext tokens, aiming to extract fine-grained knowledge specific to each pretext task from the pre-trained model.
Finally, prompting has also been considered on semantic-rich graphs, including heterogeneous and text-attributed graphs. HGPrompt Yu et al. (2023a) extends GraphPrompt to heterogeneous graphs, utilizing an additional graph template to convert a heterogeneous graph into homogeneous ones, bridging the gap between the two kinds of graph. For text-attributed graphs, G2P2 Wen and Fang (2023a) integrates graph neural networks with a language model, adopting a multi-modal alignment framework Radford et al. (2021) to pre-train the language model and graph model jointly. Alternatively, OFA Liu et al. (2024) describes subgraphs using texts, which can be subsequently fed to a pre-trained language model. Both works employ language-based templates and prompts akin to those used in NLP.
Parameter-efficient fine-tuning (PEFT). These methods only tune a selected subset of parameters from the original pre-trained model or from newly added modules. Hence, the number of parameters requiring updating is significantly reduced, which proves more efficient and feasible for few-shot learning. Prominent PEFT techniques include adapter tuning Houlsby et al. (2019) and Low-Rank Adaptation (LoRA) Hu et al. (2021). Adapter tuning Houlsby et al. (2019) inserts small neural network modules called adapters into certain layers of a pre-trained model, while LoRA Hu et al. (2021) utilizes low-rank matrices to approximate parameter updates. In both techniques, only the parameters of the new modules are updated during fine-tuning, while the original pre-trained weights remain frozen.
For pre-trained graph models, AdapterGNN Li et al. (2023a) introduces adapter-based fine-tuning into graph neural networks. This is achieved via two parallel adapters that process inputs before and after message passing. In a similar spirit, G-Adapter et al.Gui et al. (2023) refines the graph transformer architecture with adapters, and further incorporates structural information into the adapter. In a different graph-assisted application, Liu et al. Liu et al. (2023d) focus on the molecule-to-text generation task and propose molecular graph language modeling. They design a cross-modal projector aimed at bridging the gap between graph structural and textual representations, while integrating a uni-modal LoRA module into their language model.
3.4 Discussion
Unlike meta-learning approaches that require a large volume of labeled data to construct meta-learning tasks, pre-training approaches employ self-supervised pretext tasks on unlabeled data. Therefore, pre-training approaches are more effective in scenarios where labeled data are limited to novel tasks without a pre-existing set of annotated tasks. In contrast, when such an annotated base set is available, meta-learning tends to perform better as it can leverage related meta-training tasks derived from the base set. However, the dependency on related tasks suggests that meta-learning approaches cannot be easily extended to a wide array of downstream tasks, rendering them less general than pre-training approaches.
Adapting a pre-trained model to downstream tasks by fine-tuning is an intuitive strategy, but it can be ineffective due to the objective gap and overfitting to few-shot labels, or costly for a large model. Parameter-efficient adaptation strategies, including prompt tuning, adapter tuning and LoRA, present a more promising direction for few-shot learning on graphs.
4 Hybrid Approaches
Meta-learning and pre-training embody two distinct paradigms to learn prior knowledge, each with its unique advantages and disadvantages, as previously discussed. In situations where both a vast set of unlabeled data for pre-training and an extensively annotated base set for meta-learning are available, hybrid approaches that embrace both paradigms may be a natural and strategic choice for few-shot learning. Hence, we can leverage the respective strengths of each, not only exploiting the unlabeled data to learn a general task-agnostic prior via pre-training, but also drawing insights from similar meta-training tasks.
As depicted in Fig. 4(a), hybrid approaches adopt typical contrastive or generative pretext tasks to pre-train a graph encoder. Subsequently, the pre-trained model is adapted in conjunction with meta-learning, as illustrated in Fig. 4(b,c). Specifically, VNT Tan et al. (2023) employs a pre-trained model to reduce reliance on the base classes. It further proposes structure prompts, serving as virtual nodes to tailor the pre-trained node embeddings for meta-training tasks, further leveraging the insights from the base classes. However, VNT is specifically designed for node-level tasks only, falling short of handling edge- or graph-level tasks. In contrast, ProG Sun et al. (2023) reformulate the downstream node- and edge-level tasks as graph-level tasks, and proposes prompt graphs with specific nodes and structures to guide different tasks. Similar to VNT, it incorporates meta-learning to acquire initial prompts for handling different tasks. On text-attributed graphs, G2P2* Wen and Fang (2023b) extends G2P2 Wen and Fang (2023a) with conditional prompt tuning, which employs a meta-net to generate a distinct prompt vector for each node. The meta-net is trained using labeled data from base classes, facilitating knowledge transfer to novel classes.
Figure 4: An illustration of hybrid approaches that integrate meta-learning and pre-training for few-shot learning on graphs.
5 Future Directions
This section aims to identify and discuss potential opportunities for few-shot learning on graphs, leveraging the insights gained from our review of problems and techniques.
5.1 Future avenues in problem setting
Few-shot learning on large-scale graphs. Previous works predominantly concentrate on small graph structures such as molecule graphs Hu et al. (2020a), which consist of tens or hundreds of nodes only. However, many real-world applications deal with much larger graph structures, such as social networks, interaction network for recommender systems, and protein or gene networks. Challenges on these larger graphs encompass not only the immediate issues of computational efficiency and scalability, but also more finer-grained adaptation strategies to deal with potential variations among distant localities on a larger graph.
Few-shot learning on complex graphs. Most of the current works focus on simple homogeneous structures, commonly exemplified by 2D graphs, uni-modal graphs, and static graphs. However, the practical applicability of few-shot learning extends beyond these limitations to include more complex structures, such as 3D graphs Liu et al. (2021a), multi-modal graphs Li et al. (2023b), and dynamic graphs Yang et al. (2022), each with important applications in scientific, biomedical or social domains. This expansion in scope highlights the need for novel methodologies and techniques to effectively address the challenges posed by each of these intricate graph structures.
Few-shot learning on cross-domain graphs. Cross-domain few-shot learning involves transferring knowledge from one domain to another, emphasizing the generalization across diverse scenarios Yao et al. (2020). Graphs from different domains may vary in size, topology and semantic interpretation, which may call for specialized domain adaptation techniques to mitigate domain shift.
5.2 Future avenues in techniques
Improving interpretability. Enhancing interpretability is a critical future direction for few-shot learning on graphs. The black-box nature of some techniques makes it challenging to understand and interpret their results. For instance, unlike language-based prompts in NLP that can be understood by humans, most prompts designed for graphs are simply learnable vectors in the latent space. It thus becomes imperative to develop techniques that provide insights into the rationale behind predictions for few-shot tasks on graphs. Research in this direction could start with designing more interpretable prompts on graphs, explicitly explaining their relationship to node features or graph structures.
Foundation models on graphs. Foundation models pre-trained on large-scale data in NLP have shown powerful performance across various kinds of tasks in diverse application domains. Similarly, a pre-trained graph foundation model holds significant promise for diverse few-shot learning tasks on graphs. Toward graph foundation models, we aim to improve the transferability of current pre-training efforts on graphs, enabling their application in diverse domains as well as different task types. On one hand, improving transferability across domains requires not only mitigating the issue of domain shift on cross-domain graphs (see Sect. 5.1), but also bridging the gap created by heterogeneity in graph structures. This heterogeneity stems from different kinds of complex structures, including 3D, multi-modal and dynamic graphs (see Sect. 5.1). Unifying these structures is critical to the universal applicability of graph foundation models; in particular, multi-modal graphs pave the way for integrating graph foundation models with language- and vision-based models. On the other hand, there is a pressing need to enhance transferability across task types, extending beyond the conventional tasks of link prediction, node classification, and graph classification. More variety of task types, such as regression, graph editing and generation Liu et al. (2023b), can be crucial to material engineering and drug development.
6 Conclusions
In this survey, we offer a systematic yet concise review of the literature on few-shot learning on graphs. Based on predominant techniques, we categorize these studies into meta-learning approaches, pre-training approaches and hybrid approaches. For each category, we present representative works and discuss their respective advantages and drawbacks. Finally, we outline promising future directions, aiming to stimulate further exploration in this rapidly evolving field.